presently,bill is 13 and his dad is 35. what will thier ages be when the ratio of thier ages is 2:3?
Answer by lwsshak3(11628)

presently,bill is 13 and his dad is 35. what will their ages be when the ratio of thier ages is 2:3?
**
let x=years from present when ratio of their ages is 2:3
13+x/35+x=2/3
39+3x=70+2x
x=31
ans:
Bill will be 44 years of age
His dad will be 66 years of age